# Layered Insights: Generalizable Analysis of Authorial Style by Leveraging All Transformer Layers

### v1

@@ 4.2.2 Multi-domain Training
Table 2 presents the results. Overall, our MultiLUAR model benefits more than LUAR from including different training domains, achieving higher scores in almost all cases. Comparing with the single-domain results shown in Table 1, we find that introducing Reddit data consistently improves transferability, reinforcing its diverse linguistic coverage as an essential factor in cross-domain generalization. Unlike LUAR, where adding Fanfiction dampened the performance, our MultiLUAR model enhances out-of-domain generalization because it effectively leverages different linguistic granularity levels, allowing it to capture both fine-grained lexical choices and higher-level stylistic structures. This enables the model to distinguish authorial identity beyond surface-level patterns, improving performance across structurally diverse domains. This is particularly evident in models trained on Amazon $\cup$ Fanfiction; where LUAR failed, MultiLUAR has a strong performance on Reddit by integrating multiple stylistic cues. Conversely, introducing Amazon leads to mixed results, improving transferability in some cases while slightly reducing in-domain performance. Furthermore, introducing Reddit into in-domain evaluations also leads to performance improvements. Given its strong transferability, we observe enhanced in-domain performance across other datasets as well. Notably, due to MultiLUAR’s significant improvements in Fanfiction’s out-of-domain capabilities, we also see in-domain gains when paired with Fanfiction, particularly in the Amazon $\cup$ Fanfiction setting.
@@ 5 Analysis
In this section, we take an analytical look at what our model’s projection layers capture linguistically and what role they play in the final prediction in both in and out-of-domain scenarios.
@@
5.1 Learned Linguistic Representations
As mentioned, research showed that during the pre-training phase of transformer models, their layers acquire different linguistic capabilities. In this section, we look at this phenomenon in our trained MultiLUAR model to inspect whether the training process on authorship attribution task leads to different layers capturing different linguistic phenomena. In particular, we use the Holmes benchmark Waldis et al. (2024), which consists of over 200 datasets and different probing methods to evaluate the linguistic competence in transformers, including syntax, morphology, semantics, and discourse at different layers. We run the probing method of this library over paraphrase-distilroberta, the base transformer of our model, quantifying the linguistic competence of its initial embedding (layer 0) and all its 6 layers. We then compare it to MultiLUAR model that we trained on the Reddit dataset.
Figure 4: Percentage of text pairs (y-axis) having significantly high similarity score w.r.t. different layers (x-axis) of MultiLUAR trained on (a) Reddit, (b) Amazon, and (c) Fanfiction. These percentages are broken by the evaluation dataset: Reddit, Amazon, and Fanfiction
Figure 2 visualizes the linguistic competence of all layers for both models. As for RoBERTa model (Figure 2.b), we observe that for all four linguistic phenomena, layer 3 has the strongest representation rather than the final layer (a 0.5 score in Morphology and around 0.7 on Semantics, Discourse, and Syntax). While layer 6 is better at capturing semantics and morphology than layer 1 (0.6 and 0.38 compared to 0.5 and 0.3, respectively), it is worse at discourse and syntax. Since the authorship attribution task relies on comparing text similarity under different linguistic aspects rather than only semantics, these insights support the need to consider all transformer layers when solving the authorship attribution task. Furthermore, we can see that all layers are stronger at syntax, discourse, and semantics compared to morphology. Now, in comparison, looking at Figure 2.a that presents our MultiLUAR model’s layers linguistic capacity, we can observe the following. The competence of all layers is relatively stronger for morphology and syntax compared to discourse and semantics. This is an indication that syntax and morphology are more important for the task of authorship attribution than semantics and discourse, at least on the Reddit dataset that the model is trained on. Furthermore, we can see that different layers become specialized in specific linguistic phenomena. For example, layer 4 represents morphology best; layer 3 captures syntax, and layers 0 and 6 capture discourse better than other layers. One can also notice that, overall, MultiLUAR model’s linguistic capacity is lower than paraphrase-distilroberta across all layers. For example, at morphology, the highest score of MultiLUAR is 0.35 compared to 0.5 for RoBERTa. This decrease is expected since these projection layers of our model have a lower dimension size (512 compared to 768).
@@ 5.2.1 Ablation Study
The final prediction of our model is an aggregation score of the cosine similarity between all layer representations of the input pairs. Therefore, our ablation study evaluates the impact of different layers in the MultiLUAR model by systematically removing each layer from the aggregation and measuring its effect on attribution performance across three datasets: Reddit, Amazon, and Fanfic, allowing us to categorize this contribution for in and out-of-domain scenarios. Figure 3 presents the ablation results as a bar chart for each MultiLUAR model when evaluated on each of the three datasets while removing each of its layers from the predictions. Each bar reflects the change in the model’s performance in terms of R@8 after removing a specific layer. Negative bars indicate that the layer harmed the model’s performance.
MultiLUAR trained on Fanfic
@@
MultiLUAR trained on Amazon
@@
MultiLUAR trained on Reddit
@@
Overall, the findings reinforce the importance of multi-layer representations in author attribution tasks. Unlike prior work that relies solely on final-layer embeddings, our results demonstrate that early and middle layers’ embeddings contribute significantly to out-of-domain generalization, with various layers playing different importance roles based on the evaluation dataset’s style characteristics. Higher layer embeddings then capture more nuanced, domain-specific stylistic cues. The MultiLUAR approach, which aggregates information across all linguistic granularity levels, provides a more robust framework for modeling author writing style. These results highlight the limitations of single-layer methods and emphasize the necessity of incorporating multiple levels of linguistic representation for improved attribution performance.
@@ 5.2.2 Layer significance
We analyze how relevant a specific layer is for predictions on each test dataset (Reddit, Amazon, and Fanfiction). We quantify layer relevance by computing the percentage of instances in the test set (a pair of author texts) with a cosine similarity higher or lower than the average score by a given margin. Specifically, given a corresponding dataset, a MultiLUAR model, and one of its layers, we sample a set of 5k text pairs, measure the cosine similarity of their representations at all layers, and compute the percentage of cases where the z-score of the corresponding layer is higher than 1.5 for positive pairs (indicating cosine similarity 1.5 standard deviations higher than average) or lower than -1.5 for negative pairs (indicating lower cosine similarity than average). We visualize these percentages in Figure 4 for the three MultiLUAR models trained on Reddit, Amazon, and Fanfiction datasets. Looking at Figure 4, we can see an overall pattern where the final layer plays a more specific role in predicting in-domain cases while early layers of the model contribute to the predictions of the out-of-domain cases, which aligns with our ablation study presented earlier. For example, considering at MultiLUAR trained on Reddit (Figure 4.a), we observe that layers 0 and 6 are mainly relevant to the Reddit instances, while the other layers, such as layer 1, play a more prominent role in the out-of-domain scenario (Amazon and FanFic datasets). Similarly, for MultiLUAR trained on Amazon (Figure 3.b), both layers 0 and 6 are more relevant to the in-domain predictions (Amazon instances), while for out-of-domain (Reddit and Fanfiction), also early layers such as 1 to 3 contribute to the predictions. Finally, as for the MultiLUAR model trained on Fanfiction (Figure 3.c), we see a clearer pattern where layer 6 contributes mostly to the in-domain predictions while early layers (0 to 2) predict out-of-domain cases.
@@ 6 Discussion and Conclusion
In this paper, we argued that an author’s writing style presents itself at different levels of linguistic granularity. For example, two author texts can be similar (hence likely written by the same author) because they have a similar vocabulary (lexical level) or follow similar structure (discourse level). Therefore, to address the authorship attribution task more effectively, we proposed a new approach that leverages all the representations learned at different layers of the base transformer instead of using only the final layer. This feature of our approach becomes more important when evaluating models out-of-domain since utilizing multiple levels of linguistic granularity enriches the model’s representations, allowing it to perform more robustly.
We evaluated our approach against state-of-the-art baselines on three datasets. Results confirmed how our approach leads to increased performance on the task in the out-of-domain scenario while maintaining relevantly strong performance in the in-domain settings. This gap between our model and state-of-the-art becomes apparent when training is carried out on a more homogeneous dataset like Fanfic and Amazon compared to Reddit.
We then studied what linguistic representations are captured at the different layers of the base transformer (RoBERTa) that our model builds on, and we compared how these representations change in the corresponding layers of our model. Our investigation showed that while RoBERTa’s layers capture more semantics and discourse, the layers of our model capture morphology and syntax relatively better, indicating the importance of these two linguistic properties for the AA task. We further analyze the contribution of each layer in our model towards the final prediction through an ablation study. Across the three trained models, we found a trend of having layers contributing more to the out-of-domain performance than in-domain, such as layers 0 and 1 in MultiLUAR trained on Reddit. Latter layers of all models were found to be essential to the in-domain performance. Overall, our findings demonstrate the importance of leveraging all transformer layers when addressing the authorship attribution task to achieve robust performance in out-of-domain scenarios.
@@ 7 Limitations
First, we applied our approach to only one authorship attribution model, which is LUAR Rivera-Soto et al. (2021). Although we tested our hypothesis on three different datasets and showed a consistent behavior, one can potentially demonstrate the robustness of our method by applying it to other transformer-based AA models.
Second, in our analysis of the linguistic competence of each of the layers of our MultiLUAR model, we rely on some probing methods. However, these might not be very reliable tools.
@@
Besides, state-of-the-art systems for authorship attribution might suffer from generating false positives, for example, scenarios where a post is wrongly attributed to an individual in a criminal case. Such false positives become more frequent in out-of-domain scenarios. We believe that the research in this paper is a step towards building more robust authorship attribution systems.
Acknowledgements
This research is supported in part by the Office of the Director of National Intelligence (ODNI), Intelligence Advanced Research Projects Activity (IARPA), via the HIATUS Program contract #2022-22072200005. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of ODNI, IARPA, or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein
@@ Appendix A Experimental Setup
For our experiments, we utilize the RoBERTa model (paraphrase-distilroberta) Sanh et al. (2020) of 82.5 M parameters size, including seven additional linear layers, as the backbone for both the MultiLUAR approach and the baseline.
@@
Training is conducted using an NVIDIA A100 GPU (40GB memory) to handle the computational demands of large-scale transformer-based models. Our training spans three datasets : Reddit, Fanfiction, and Amazon reviews with a fixed training schedule of 20 epochs per dataset. Specifically, training on the Reddit dataset requires approximately 1 hour per epoch, while Fanfiction and Amazon reviews take 11 minutes and 5 minutes per epoch, respectively.
We do not perform any hyperparameter tuning, opting instead to use the default parameters from the LUAR implementation. Though we experimented with training beyond 20 epochs with some minute performance gains, we adhere to the original LUAR training setup to ensure reproducibility of prior results.
Our training pipeline uses PyTorch Lightning (1.8.5.post0), PyTorch (1.13.1), Transformers (4.14.1), and Scikit-learn (1.2.0), which facilitate efficient model training, evaluation, and data handling.
Datasets Content.
@@
Appendix B Ablation Study
B.1 Significance of Layer 0
Our model, MultiLUAR, consists of seven projection layers, where Layer 0 represents the projection of the initial token embeddings from the distillroberta transformer, while Layers 1 to 6 correspond to the projection of its six hidden layers. To demonstrate that, although Layer 0 is not part of the model architecture, it plays a significant role in authorship attribution, we train and evaluate all models on the three datasets without adding a linear projection to Layer 0. The embedding is not incorporated into the loss calculation and similarity metric computation. The results in Table 3 compare the performance of MultiLUAR with and without Layer 0 across different evaluation datasets. We observe that including Layer 0 generally improves performance in out-of-domain evaluation, particularly on the Amazon and Fanfiction datasets, where it enhances R@8 and MRR scores. This suggests that initial token embeddings contain valuable stylistic information that contributes to generalization across domains. However, for in-domain evaluation on Reddit, the model without Layer 0 slightly outperforms the full model as shown in section 5.2.1, indicating that deeper layers are more influential in capturing dataset-specific stylistic patterns. These findings reinforce the importance of leveraging multi-layer representations for robust authorship attribution.

### v2

@@ 5.2 Multi-domain Training
In this experiment, we train LUAR and our LIGHT-LUAR on two domains and evaluate them on the third one to examine whether exposure to multiple domains enhances robustness. To this end, we train models on domain pairs using mini-batches of 256 that are randomly selected from document collections, ensuring equal representation from both domains. Since authors are disjoint across domains, shared authorship occurs only within the same domain. Table 2 presents the results. Overall, our LIGHT-LUAR model benefits more than LUAR from including different training domains, achieving higher scores in almost all cases. Compared with the single-domain results shown in Table 1, we find that introducing Reddit data consistently improves transferability, reinforcing its diverse linguistic coverage as an essential factor in cross-domain generalization. Unlike LUAR, where adding Fanfiction dampened the performance, our LIGHT-LUAR model enhances out-of-domain generalization because it leverages different linguistic granularities. This is evident in models trained on Amazon $\cup$ Fanfiction; where LUAR failed, LIGHT-LUAR has strong performance on Reddit by integrating multiple stylistic cues.
@@ 6 Analysis
In this section, we look at what our model’s projection layers capture linguistically and what role they play in the final prediction in both in and out-of-domain scenarios. All analyses are run when our approach is applied to the LUAR (LIGHT-LUAR) model.
6.1 Learned Linguistic Representations
By looking at what linguistic capacities are acquired at the different layers, the goal is to demonstrate the rich linguistic representations captured in our model compared to previous state-of-the-art models. In particular, we use the Holmes benchmark Waldis et al. (2024), which consists of over 200 datasets and different probing methods to evaluate the linguistic competence of representations, including syntax, morphology, semantics, and discourse. We run the probing method of this library over each output layer of our LIGHT-LUAR model and on the single output layer of LUAR (both are trained on the Reddit dataset).
Figure 2 shows a heatmap of the linguistic competence of all layers of the two models. We can observe that LUAR’s output layer (first column) has a huge skew towards morphology compared to other linguistic phenomena, indicating the importance of morphological representation in texts for the AA task. In contrast, our LIGHT-LUAR model has a more balanced representation of all the linguistic phenomena, with the highest being morphology captured at Layer 4 (L 4). Aside from Morphology, our LIGHT-LUAR represents each linguistic phenomenon at one of its layers better than LUAR’s final output layer. Aligned with previous literature, Syntax is best captured at the intermediate layers (L 2-3), while discourse is at the later layers (L 3-6), as well as L 0, potentially, capturing discourse markers. These observations indicate the need to model authorship attribution at different layers to benefit from the diverse captured linguistic representation, rather than considering only the final output layer.
@@ 6.2.1 Ablation Study
The final prediction of our model is an aggregation score of the cosine similarity between all layer representations of the input pairs. Therefore, our ablation study evaluates the impact of different layers in the LIGHT-LUAR model by systematically removing each layer from the aggregation and measuring its effect on attribution performance across three datasets: Reddit, Amazon, and Fanfic, allowing us to categorize this contribution for in and out-of-domain scenarios. Figure 3 presents the ablation results as a bar chart for each LIGHT-LUAR model when evaluated on each of the three datasets while removing each of its layers from the predictions. Each bar reflects the change in the model’s performance in terms of R@8 after removing a specific layer. Negative bars indicate that the layer harmed the model’s performance.
LIGHT-LUAR trained on Fanfic
@@
LIGHT-LUAR trained on Amazon
@@
LIGHT-LUAR trained on Reddit
@@
Figure 4: Percentage of text pairs (y-axis) having significantly high similarity score w.r.t. different layers (x-axis) of LIGHT-LUAR trained on (a) Reddit, (b) Amazon, and (c) Fanfiction. These percentages are broken by the evaluation dataset: Reddit, Amazon, and Fanfiction
Overall, the findings reinforce the importance of multi-layer representations in author attribution tasks. Unlike prior work that relies solely on final-layer embeddings, our results demonstrate that early and middle layers’ embeddings contribute significantly to out-of-domain generalization, with various layers playing different importance roles based on the evaluation dataset’s style characteristics. Higher layer embeddings then capture more nuanced, domain-specific stylistic cues. The LIGHT-LUAR approach, which aggregates information across all linguistic granularity levels, provides a more robust framework for modeling author writing style. These results highlight the limitations of single-layer methods and emphasize the necessity of incorporating multiple levels of linguistic representation for improved attribution performance.
@@ 6.2.2 Layer significance
We analyze how relevant a specific layer is for predictions on each test dataset (Reddit, Amazon, and Fanfiction). We quantify layer relevance by computing the percentage of instances in the test set (a pair of author texts) with a cosine similarity higher or lower than the average score by a given margin. Specifically, given a corresponding dataset, a LIGHT-LUAR model, and one of its layers, we sample a set of 5k text pairs, measure the cosine similarity of their representations at all layers, and compute the percentage of cases where the z-score of the corresponding layer is higher than 1.5 for positive pairs (indicating cosine similarity 1.5 standard deviations higher than average) or lower than -1.5 for negative pairs (indicating lower cosine similarity than average). We visualize these percentages in Figure 4 for the three LIGHT-LUAR models trained on Reddit, Amazon, and Fanfiction datasets. Looking at Figure 4, we can see an overall pattern where the final layer plays a more specific role in predicting in-domain cases while early layers of the model contribute to the predictions of the out-of-domain cases, which aligns with our ablation study presented earlier. For example, considering at LIGHT-LUAR trained on Reddit (Figure 4.a), we observe that layers 0 and 6 are mainly relevant to the Reddit instances, while the other layers, such as layer 1, play a more prominent role in the out-of-domain scenario (Amazon and FanFic datasets). Similarly, for LIGHT-LUAR trained on Amazon (Figure 3.b), both layers 0 and 6 are more relevant to the in-domain predictions (Amazon instances), while for out-of-domain (Reddit and Fanfiction), also early layers such as 1 to 3 contribute to the predictions. Finally, as for the LIGHT-LUAR model trained on Fanfiction (Figure 3.c), we see a clearer pattern where layer 6 contributes mostly to the in-domain predictions while early layers (0 to 2) predict out-of-domain cases.
@@ 7 Discussion and Conclusion
Starting from the observation that an author’s writing style presents itself at different granularities, we proposed a new approach that leverages all the representations learned at different layers of the base transformer to model the AA task more effectively. This feature of our approach becomes more important in out-of-domain scenarios since utilizing multiple levels of linguistic granularity enriches the model’s representations, allowing it to perform more robustly. We argue that our approach also allows better interpretability of predictions since authors’ similarities can be analyzed on different linguistic levels, allowing a more fine-grained understanding of the prediction.
We evaluated our approach by applying it to two popular AA models and evaluated it on three datasets. Results confirmed how our approach leads to increased performance on the task in the out-of-domain scenario while maintaining relatively strong performance in the in-domain settings. We then studied what linguistic representations were captured at the different layers of our model, demonstrating the reason behind its reliable and robust performance. We further analyze the contribution of each layer in our model towards the final prediction through an ablation study. Across the three trained models, we found a trend of having layers contributing more to the out-of-domain performance than in-domain. Later layers of all models were found to be essential to the in-domain performance.
@@ 8 Limitations
First, our model architecture implies adding extra parameters to the original AA structure that require more training time and computational resources, especially when the number of layers of the base transformer increases. Future experiments could consider whether all layers need to be projected in order to achieve the best performance. Second, although we verified our hypothesis on two different AA models with two different base transformers, we focused the presentation of our analysis only on the LUAR model due to space limitations. Third, we rely on probing methods to analyze the linguistic competence of our projected layers. However, these might not be very reliable tools, and they have their own limitations.
@@
Besides, state-of-the-art systems for authorship attribution might suffer from generating false positives, for example, scenarios where a post is wrongly attributed to an individual in a criminal case. Such false positives become more frequent in out-of-domain scenarios. We believe that the research in this paper is a step towards building more robust authorship attribution systems that allow fine-grained analysis of what specific linguistic similarities have been detected, granting more system interpretability.
@@ A.1 LUAR
For our experiments, we utilize the RoBERTa model (paraphrase-distilroberta) Sanh et al. (2020) with 82.5M parameters and seven additional linear layers as the backbone for both the LIGHT-LUAR approach and the baseline. We additionally experiment with ModernBERT (ModernBert-base)Warner et al. (2024), a larger architecture comprising 149M parameters and 23 additional linear layers including embedding layer, which enables deeper stylistic representation and broader contextual modeling across layers.
@@
Training is conducted using an NVIDIA A100 GPU (40GB memory) to handle the computational demands of large-scale transformer-based models. For RoBERTa-based models, training spans three datasets - Reddit, Fanfiction, and Amazon Reviews, with a fixed schedule of 20 epochs per dataset. Specifically, training on the Reddit dataset requires approximately 1 hour per epoch, while Fanfiction and Amazon take 11 minutes and 5 minutes per epoch, respectively. We do not perform any hyperparameter tuning, opting instead to use the default parameters from the LUAR implementation. Although training beyond 20 epochs yielded minor performance gains, we adhere to the original LUAR setup to ensure reproducibility.
For ModernBERT, training is conducted on the same three datasets with a fixed schedule of 30 epochs per dataset. Due to the model’s increased depth and wider context window, training is significantly slower: the Reddit dataset requires approximately 4 hours per epoch, while Fanfiction and Amazon take 1 hour 10 minutes and 12 minutes per epoch, respectively.
Our training pipeline utilizes PyTorch Lightning (2.5.1.post0), PyTorch (2.5.1+cu121), Transformers (4.51.3), and Scikit-learn (1.6.1), which together facilitate efficient model training, evaluation, and data handling. Since ModernBERT requires Transformers version $\geq$ 4.49.0, we recommend installing PyTorch first, followed by a compatible, upgraded version of Transformers.
A.2 Wegmann
Model.
We evaluate two backbone models: RoBERTa (125M parameters) and ModernBERT (149M parameters), both initialized with pretrained HuggingFace weights. Following the Sentence-BERT framework Reimers and Gurevych (2019), we implement two variants: a single-layer model, using only the final-layer [CLS] token for embedding; and a proposed all-layer model that incorporates structural modifications to support fine-grained supervision.
The all-layer model uses a custom transformer module (TransformerAllLayers) that forces the underlying transformer to output all intermediate hidden states by enabling output_hidden_states=True. These hidden states (a tuple of shape (L, B, T, H) for $L$ layers, batch $B$, tokens $T$, and hidden size $H$) are extracted from all hidden layers and passed through the training pipeline for supervision.
To exploit these, we introduce a MultiLayerTripletLoss objective. For each training triplet (anchor, positive, negative), we extract the [CLS] token from every layer and compute layer-wise cosine distances between anchor-positive and anchor-negative pairs. The loss is defined as $\frac{1}{L}\sum_{\ell=1}^{L}\max\left(0,d_{\cos}(a^{\ell},p^{\ell})-d_{\cos}(a^{\ell},n^{\ell})+m\right)$, where $a^{\ell},p^{\ell}$, and $n^{\ell}$ are the [CLS] embeddings from layer $\ell$, and $m$ is the margin (set to 0.5). This design encourages style-sensitive representations to emerge throughout the network depth, rather than only in the final layer.
Training.
Training is performed on 59k Reddit-MUD triplets and 210k Wegmann conversation triplets. We use a 4×NVIDIA A100 GPU cluster (80GB) for efficient multi-GPU training. Epoch runtimes vary by model variant and dataset (Reddit and Wegmann data respectively):
  • All-layer: 28 min, 80 min
  • Single-layer: 20 min, 52 min
We use early stopping with a $10^{-4}$ threshold and patience of 3 epochs. Training typically converges within 9 epochs. Tokenization edge cases are handled robustly, with input sequences truncated to 512 wordpieces using the HuggingFace tokenizer. Our implementation stack includes PyTorch (2.7.0), Transformers (4.51.3), SentenceTransformers (4.1.0), and Scikit-learn (1.6.1).
Evaluation.
We follow LUAR’s protocol by ranking each author’s utterance embeddings against a candidate set. However, unlike LUAR, which uses classification logits, Wegmann evaluates directly on sentence embeddings from the trained model (either single-layer or all-layer). Embedding similarities are computed via cosine similarity.
To reduce inference time (during evaluation only), we enable torch.bfloat16 precision for both models on A100 hardware. We verified numeric stability by comparing evaluation scores with full-precision outputs and found no degradation in accuracy. This optimization yields an effective $\sim$10× speedup.
Evaluation runtimes (ModernBERT, RoBERTa):
  • Reddit: 24 min, 14 min
  • Fanfiction: 6 min, 4 min
  • Amazon: 120 min, 40 min
A.3 Datasets
@@
Appendix B ModernBERT Results
We evaluated our methodology using ModernBERT as the foundational transformer model. The outcomes demonstrate consistent and significant improvements over the RoBERTa-based models discussed previously, attributable to ModernBERT’s increased context length and greater model depth. Results are shown in Table 3
For in-domain evaluation, the performance patterns align closely with previous observations using RoBERTa. Incorporating multiple transformer layers with ModernBERT-based LUAR and Wegmann approaches typically results in comparable or marginally reduced performance relative to their single-layer counterparts. This indicates that within the same domain, the deeper final-layer representations from ModernBERT already encapsulate extensive stylistic and content information, minimizing the benefits derived from incorporating intermediate layers. Nonetheless, we note distinct improvements in scenarios involving structured writing styles, such as the Amazon dataset, where leveraging multiple layers evidently captures subtle stylistic features more effectively than using the final layer alone.
In the out-of-domain evaluation setting, however, our multi-layer approach with ModernBERT substantially enhances model generalizability and sets new benchmarks compared to the previously used RoBERTa-based models. The expanded contextual understanding and richer intermediate representations provided by ModernBERT significantly improve cross-domain adaptability. For instance, ModernBERT-based LIGHT-LUAR trained on structured datasets like Amazon demonstrates notably superior adaptability when evaluated on informal, conversational texts like Reddit and Fanfiction, capturing nuanced stylistic features that generalize across diverse domains. Similarly, when models trained on conversational datasets such as Reddit are evaluated on structured writing domains (Amazon), the ModernBERT variants consistently outperform previous baselines.
Most notably, on the challenging Fanfiction dataset, known for its diverse and expressive language usage, our ModernBERT-based LIGHT-LUAR and LIGHT-Wegmann models exhibit remarkable improvements, highlighting their ability to generalize across significantly different linguistic styles. These improvements underscore the value of leveraging multiple transformer layers in ModernBERT, particularly for enhancing performance in complex, stylistically rich domains.
Training Dataset